# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ 4.1 Barrier Certificates Based Invariance Verification
It is worth remarking here that we do not extend the k-Inductive Barrier certificates proposed in [3] for addressing the problem in this paper. A set of sufficient conditions, which is similar to the one in Proposition 2 can be easily obtained based on k-Inductive barrier certificates. However, the gain of sufficient conditions being analogous to the ones in Proposition 3 and 4.4 should be carefully treated and will be considered in the future work.
@@ 4.2 Equations Relaxation Based Invariance Verification
In this subsection, we present the second set of optimizations for addressing the safe invariance verification problem in Definition 2. We begin by introducing an equation that characterizes the liveness probability of staying the safe set $\mathcal{X}$. This equation is adapted from (3), where the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ is regarded as the target set $\mathcal{X}_{r}$ in (3). We then propose two sufficient conditions for certifying lower and upper bounds of the liveness probability by relaxing the derived equation.
@@
Given a safe set $\mathcal{X}$, if there exist a function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$  [^2] Comparing with Theorem 1, the explicit requirement that $v(\bm{x})$ is bounded is abandoned here. If following the proof of Theorem 1 in [29], one can find that this requirement is not necessary. and a bounded function $w(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$ such that for $\bm{x}\in\widehat{\mathcal{X}}$,
$\begin{cases}v(\bm{x})=\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]\\ v(\bm{x})=1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})+\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x})\end{cases},$ (13)
then
$\begin{split}&v(\bm{x}_{0})=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\\ &=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\\ &=\lim_{i\rightarrow\infty}\frac{\mathbb{E}^{\infty}[\sum_{j=0}^{i-1}1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j))]}{i}.\end{split}$
Thereby, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=1-v(\bm{x}_{0})$.
Proof 4.12.
Like Proposition 1, two sufficient conditions can be obtained for certifying lower and upper bounds of the liveness probability via directly relaxing equation (13), respectively.
@@
$\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (14)
then
$\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\leq v(\bm{x}_{0})\leq 1-\epsilon_{1}.$
Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$.
Proof 4.14.
According to Proposition 4.13, a lower bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op3}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),w(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ \epsilon_{1}\geq 0\\ w(\bm{x})\text{~{}is bounded over~{}}\widehat{\mathcal{X}}\end{cases}.\end{split}$
Lemma 4.15.
@@
$\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]\leq v(\bm{x}_{0})<-\delta,\forall k\in\mathbb{N}.$ (15)
Also, since $v(\bm{x})\geq 1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})+\mathbb{E}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x}),\forall\bm{x}\in\widehat{\mathcal{X}}$, we have
$v(\bm{x})\geq\mathbb{E}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x}),\forall\bm{x}\in\widehat{\mathcal{X}}.$
@@
$\begin{split}(j+1)v(\bm{x}_{0})&\geq\sum_{k=0}^{j}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]\\ &\geq\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1))]-w(\bm{x}_{0}),\forall j\in\mathbb{N}.\end{split}$
@@
$\begin{split}\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})\leq&\sum_{k=0}^{j}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]\\ <&-\delta(j+1),\ \forall j\in\mathbb{N},\end{split}$
@@
$\frac{\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})}{j+1}<-\delta,\forall j\in\mathbb{N}.$ (16)
Since
$\lim_{j\rightarrow\infty}\frac{\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})}{j}=0,$
which contradicts (16). Therefore, we can conclude that if $v(\bm{x})$ satisfies (14), $v(\bm{x})\geq 0$ holds for $\bm{x}\in\widehat{\mathcal{X}}$. Thus, conditions (9) and (14) are equivalent.
Proposition 4.17.
@@
$\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (17)
then
$\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq v(\bm{x}_{0})\geq 1-\epsilon_{2}.$
Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq\epsilon_{2}$.
Proof 4.18.
@@
According to Proposition 4.17, an upper bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op4}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),w(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ \epsilon_{2}\geq 0\\ w(\bm{x})\text{~{}is bounded over~{}}\widehat{\mathcal{X}}\end{cases}.\end{split}$
@@ 5.1 Problem Statement
[15] Let $\bm{X}^{\bm{x}}(t,\bm{w})$ be a time-homogeneous Itô diffusion given by SDE(18) with initial state $x\in\mathbb{R}^{n}$. The infinitesimal generator $\mathcal{A}$ of $\bm{X}^{\bm{x}}(t,\bm{w})$ is defined by
$\begin{split}\mathcal{A}f(\bm{x})&=\lim_{t\rightarrow 0}\frac{E\left[f\left(\bm{X}^{\bm{x}}(t,\bm{w})\right)\right]-f(\bm{x})}{t}\\ &=\frac{\partial f(\bm{x})}{\partial\bm{x}}\bm{b}(\bm{x})+\frac{1}{2}\text{tr}(\bm{\sigma}(\bm{x})^{\top}\frac{\partial^{2}f(\bm{x})}{\partial\bm{x}^{2}}\bm{\sigma}(\bm{x}))\end{split}$
@@
Theorem 5.20 (Dynkin’s formula, [15]).
@@
Similar to Definition 2, we define the safe probabilistic invariance verification problem for continuous-time systems.
@@
Given a safe set $\mathcal{X}\subset\mathbb{R}^{n}$, which is bounded and open, and an initial set $\mathcal{X}_{0}\subseteq\mathcal{X}$, the safe probabilistic invariance verification for system (18) is to compute lower and upper bounds, denoted by $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ respectively, for the liveness probability that the system, starting from any state in $\mathcal{X}_{0}$, will remain inside the safe set $\mathcal{X}$ for all time, i.e., to compute $\epsilon_{1}$ and $\epsilon_{2}$ such that
$\epsilon_{1}\leq\mathbb{P}(\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0})\leq\epsilon_{2}.$ (19)
@@ 5.2 Reachability Probability Characterization in [31]
$\begin{cases}\widetilde{\mathcal{A}}v(\bm{x})=0\\ v(\bm{x})=1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})\end{cases},$ (20)
where the infinitesimal generator $\widetilde{\mathcal{A}}$ is
$\widetilde{\mathcal{A}}f(\bm{x})=\begin{cases}\mathcal{A}f(\bm{x})&\text{if}\ \bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ 0&\text{if}\ \bm{x}\in\partial\mathcal{X}\cup\mathcal{X}_{r}\end{cases}.\$
Then
@@
where
$\tau^{x_{0}}(\bm{w})=\inf\{\,t\mid\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\vee\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}_{r}\,\}.$
is the first time of exit of $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ from the open set $\mathcal{X}\setminus\mathcal{X}_{r}$.
By relaxing (20), the sufficient conditions for certifying lower and upper bounds of the probability $\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}$ are shown in Proposition 5.23 and 5.24, respectively.
Proposition 5.23.
@@
$\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \widetilde{\mathcal{A}}v(\bm{x})\geq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\leq 1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})&\forall\bm{x}\in\overline{\mathcal{X}}\end{cases},$ (21)
@@
$\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ \mathcal{A}u(\bm{x})-v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ -v(\bm{x})\geq 0&\forall\bm{x}\in\partial{\mathcal{X}}\\ 1-v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}_{r}\end{cases},$ (22)
then $\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}=\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\widetilde{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$ holds.
Proposition 5.24.
@@
$\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \widetilde{\mathcal{A}}v(\bm{x})\leq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})&\forall\bm{x}\in\overline{\mathcal{X}}\end{cases},$ (23)

### v2

@@ 4.1 Doob’s Nonnegative Supermartingale Inequality Based Invariance Verification
It is worth remarking here that we do not extend the k-Inductive Barrier certificates proposed in [3] for addressing the problem in this paper. A set of sufficient conditions, which is similar to the one in Proposition 4.1 can be easily obtained based on k-Inductive barrier certificates. However, the gain of sufficient conditions being analogous to the ones in Theorem 4.3 and 4.6 should be carefully treated and will be considered in the future work.
@@ 4.2 Equations Relaxation Based Invariance Verification
In this subsection, we present the second set of optimizations for addressing the safe invariance verification problem in Definition 3.2. We begin by introducing an equation that characterizes the liveness probability of staying the safe set $\mathcal{X}$. This equation is adapted from (3), where the unsafe set $\widehat{\mathcal{X}}\setminus\mathcal{X}$ is regarded as the target set $\mathcal{X}_{r}$ in (3). We then propose two sufficient conditions for certifying lower and upper bounds of the liveness probabilities by relaxing the derived equation.
Lemma 4.13.
Given a safe set $\mathcal{X}$, if there exist a function $v(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$  [^2] Comparing with Proposition 3.4, the explicit requirement that $v(\bm{x})$ is bounded is abandoned here. If following the proof of Proposition 3.4 in [30], one can find that this requirement is not necessary. and a bounded function $w(\bm{x})\colon\widehat{\mathcal{X}}\rightarrow\mathbb{R}$ such that for $\bm{x}\in\widehat{\mathcal{X}}$,
(13) $\begin{cases}v(\bm{x})=\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]\\ v(\bm{x})=1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})+\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x})\end{cases},$
then for $\bm{x}_{0}\in\mathcal{X}$,
$\begin{split}&v(\bm{x}_{0})=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}\\ &=\lim_{i\rightarrow\infty}\frac{\mathbb{E}^{\infty}[\sum_{j=0}^{i-1}1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j))]}{i}.\end{split}$
Thereby, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}=1-v(\bm{x}_{0})$, $\forall\bm{x}_{0}\in\mathcal{X}$.
Proof 4.14.
Like Corollary 3.5, two sufficient conditions can be obtained for certifying lower and upper bounds of the liveness probabilities via directly relaxing equation (13), respectively.
@@
(14) $\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}\leq v(\bm{x}_{0})\leq 1-\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 4.16.
According to Theorem 4.15, a lower bound of the liveness probabilities can be computed via solving Op3 (shown in Appendix B).
@@
(15) $\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]\leq v(\bm{x}_{0})<-\delta,\forall k\in\mathbb{N}.$
Also, since $v(\bm{x})\geq 1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})+\mathbb{E}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x})$ for $\forall\bm{x}\in\widehat{\mathcal{X}}$, we have $v(\bm{x})\geq\mathbb{E}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x}),\forall\bm{x}\in\widehat{\mathcal{X}}.$
@@
$\begin{split}(j+1)v(\bm{x}_{0})&\geq\sum_{k=0}^{j}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1))]-w(\bm{x}_{0}),\forall j\in\mathbb{N}.\end{split}$
@@
$\begin{split}\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})\leq&\sum_{k=0}^{j}\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k))]<-\delta(j+1),\ \forall j\in\mathbb{N},\end{split}$
@@
(16) $\frac{\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})}{j+1}<-\delta,\forall j\in\mathbb{N}.$
which contradicts
$\lim_{j\rightarrow\infty}\frac{\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j+1)]-w(\bm{x}_{0})}{j}=0.$
Therefore, we can conclude that if $v(\bm{x})$ satisfies (14), $v(\bm{x})\geq 0$ holds for $\bm{x}\in\widehat{\mathcal{X}}$. Thus, conditions (9) and (14) are equivalent.
Theorem 4.19.
@@
(17) $\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0},\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\widehat{\mathcal{X}}\setminus\mathcal{X}\big{)}\geq v(\bm{x}_{0})\geq 1-\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Consequently, $\mathbb{P}^{\infty}\big{(}\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}\big{)}\leq\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 4.20.
@@
According to Theorem 4.19, an upper bound of the liveness probabilities can be computed via solving Op4 (shown in Appendix B).
@@ 5.1 Problem Statement
[16] Let $\bm{X}^{\bm{x}}(t,\bm{w})$ be a time-homogeneous Itô diffusion given by SDE(18) with initial state $x\in\mathbb{R}^{n}$. The infinitesimal generator $\mathcal{A}$ of $\bm{X}^{\bm{x}}(t,\bm{w})$ is
$\begin{split}\mathcal{A}f(\bm{x})&=\lim_{t\rightarrow 0}\frac{E\left[f\left(\bm{X}^{\bm{x}}(t,\bm{w})\right)\right]-f(\bm{x})}{t}=\frac{\partial f(\bm{x})}{\partial\bm{x}}\bm{b}(\bm{x})+\frac{1}{2}\text{tr}(\bm{\sigma}(\bm{x})^{\top}\frac{\partial^{2}f(\bm{x})}{\partial\bm{x}^{2}}\bm{\sigma}(\bm{x}))\end{split}$
@@
Lemma 5.2 (Dynkin’s formula, [16]).
@@
Similar to Definition 3.2, we define the safe probabilistic invariance verification problem for continuous-time systems.
@@
Given a safe set $\mathcal{X}\subset\mathbb{R}^{n}$, which is bounded and open, and an initial set $\mathcal{X}_{0}\subseteq\mathcal{X}$, the safe probabilistic invariance verification for system (18) is to compute lower and upper bounds, denoted by $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ respectively, for the liveness probabilities that the system, starting from $\mathcal{X}_{0}$, will remain inside the safe set $\mathcal{X}$ for all time, i.e., to compute $\epsilon_{1}$ and $\epsilon_{2}$ such that
(19) $\epsilon_{1}\leq\mathbb{P}(\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X})\leq\epsilon_{2},\forall\bm{x}_{0}\in\mathcal{X}_{0}.$
@@ 5.2 Reachability Probability Characterization in [32]
(20) $\begin{cases}\widetilde{\mathcal{A}}v(\bm{x})=0\\ v(\bm{x})=1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})\end{cases},$
where
$\widetilde{\mathcal{A}}v(\bm{x})=\begin{cases}\mathcal{A}v(\bm{x})&\text{if}\ \bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ 0&\text{if}\ \bm{x}\in\partial\mathcal{X}\cup\mathcal{X}_{r}\end{cases},\widetilde{\mathcal{A}}u(\bm{x})=\begin{cases}\mathcal{A}u(\bm{x})&\text{if}\ \bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ 0&\text{if}\ \bm{x}\in\partial\mathcal{X}\cup\mathcal{X}_{r}\end{cases}.\\ \$
Then for $\bm{x}_{0}\in\mathcal{X}$,
@@
where $\tau^{x_{0}}(\bm{w})=\inf\{\,t\mid\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\vee\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}_{r}\,\}$ is the first time of exit of $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ from the open set $\mathcal{X}\setminus\mathcal{X}_{r}$.
By relaxing (20), the sufficient conditions for certifying lower and upper bounds of $\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)},\forall\bm{x}_{0}\in\mathcal{X}_{0}$ are shown in Corollary 5.5 and 5.6, respectively.
Corollary 5.5.
@@
(21) $\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \widetilde{\mathcal{A}}v(\bm{x})\geq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\leq 1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})&\forall\bm{x}\in\overline{\mathcal{X}}\end{cases},$
@@
(22) $\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ \mathcal{A}u(\bm{x})-v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ -v(\bm{x})\geq 0&\forall\bm{x}\in\partial{\mathcal{X}}\\ 1-v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}_{r}\end{cases},$
then
$\begin{split}&\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}\\ &=\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\widetilde{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\big{)}\geq\epsilon_{1},\forall\bm{x}_{0}\in\mathcal{X}_{0}.\end{split}$
Corollary 5.6.
@@
(23) $\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \widetilde{\mathcal{A}}v(\bm{x})\leq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1_{\mathcal{X}_{r}}(\bm{x})+\widetilde{\mathcal{A}}u(\bm{x})&\forall\bm{x}\in\overline{\mathcal{X}}\end{cases},$